# Advantages of Neural Population Coding for Deep Learning

### v3

@@ 1 Introduction
In the mammalian cortex, many variables, e.g., object orientation [13] and movement direction [6], have been found to be encoded by populations of neurons. In a population code, each neuron responds maximally to its preferred value and partially to other values of the encoded variable, with activation levels shaped by tuning curves such as Gaussian or cosine functions of the distance from the neuron’s preferred value. The activity of the group of neurons resembles a probability distribution of the encoded variable [18].
@@
Commonly used artificial neural networks, such as convolutional neural networks (CNNs) and multi-layer perceptrons (MLPs), encode information also by groups of neurons, particularly in their intermediate layers [8]. However, in the output layer, information is typically mapped onto different representations. For classification tasks, outputs are commonly represented as one-hot vectors, using one neuron for each classification label. For prediction tasks, output neurons typically correspond to the variables of interest, such as the position and orientation of an object in an image.
@@ 2.1 Single-Variable Output
For the weights and bias to converge, that following equality must hold,
@@
Since any $b$ value will fulfill this equality, the weights are defined only up to a constant bias term that shifts all weight values. The actual value of $b$ will depend on network’s initialization.
Next, we probe the network’s robustness to noise in the input. For this test, we perturb the input ${\bf x}$ by ${\bf d}$, so that ${\bf x}+{\bf d}$ is the new input. Moreover, let $d_{j}=a\delta_{kj}$ for a given perturbed neuron $k$, and we use the same ${\bf x}$ and target, $i/n$, as above. As a result, the output equals,
@@ 2.2 Population-Code Output
$w_{ij}^{t+1}=w_{ij}^{t}+\eta\left(\hat{y}_{i}-\sum_{k}w_{ik}^{t}x_{k}-b_{i}\right)x_{j}\,.$ (12)
For our theoretical analysis, we first use a simplified population code that resembles a one-hot vector, i.e., only one neuron, $x_{i}$, with preferred value $i/n$ gets active, and all other neuronal activations are zero. Later, we will show the difference between population code and one-hot vector.
@@
Using the same training input as above, the weight update simplifies to
$w_{ij}^{t+1}=w_{ij}^{t}+\eta\left(\delta_{ij}-w_{ij}^{t}-b_{i}\right)\,.$ (13)
@@
Again, we perturb the input by $d_{j}=a\delta_{kj}$. At the corresponding output neuron $k$, using (11) and (14), the activation will be
$y_{k}=(1-b_{k})a+b_{k}\,.$ (16)
Here, we used $k\neq i$ if $i$ is the target index of the real signal since we probe the condition for causing a failure. For causing a failure, a necessary condition is $y_{k}>1$ to shift the maximum away from the target $i$, where $y_{i}=1$. According to (16), $y_{k}>1$ is equivalent to $a>1$ if $b_{k}<1$. The bias will indeed be smaller 1 if we assume that $w_{ij}$ and $b_{i}$ are initialized to zero mean and zero, respectively (a common setting) because (13) will push $w_{ii}$ to positive values and $b_{i}=1-w_{ii}$ according to (14).
For $a<1$, the failure rate remains 0%. Thus, the population code leads to a trained network that is more robust to noise compared to one with single variable output.
For an actual population code, the activation values follow a tuning curve rather than being binary. So, the allowable threshold for $a$ is slightly smaller due to the non-zero activation of off-target neurons. For Gaussian tuning, a failure occurs if the max of the sum of the Gaussians from real signal and perturbation is shifted by at least $2/n$ compared to the target value. This condition necessitates that the activation at target is smaller than this max (tuning width: $\sigma$),
@@
where $4$ is the worst-case difference between $i$ and $k$ that could result in errors above the tolerance of $1.5/n$ for $a<1$. The right-hand side is maximized at $x=2/n$ if limiting $x$ to discrete values in steps of $1/n$. The maximum cannot be at $x>2/n$ if $a<1$ because the target Gaussian has more weight, and $x=1/n$ would be below the error tolerance.
@@ 3.1 Methods
We used various numbers of stacked linear layers of size $n=20$, except for the output layer for the single-variable approach, which consisted of one neuron. In the architecture, each hidden layer was followed by a leaky ReLU function. For the population-code approach, a sigmoid function was applied to the output, and we used Gaussian tuning curves with a width of $\sigma=0.1$.
@@
Figure 1: Robustness to input perturbations comparing single-variable, one-hot vector, and population-code outputs. Here, the networks were trained without noise. The dashed thin red line in the first panel shows the theoretical bound of $a$ below which the population-code has zero errors. Experimental results show means $\pm$ SD indicated by shaded areas (100 simulation runs).
@@ 3.2 Results on Noise Robustness
Figure 2: Robustness to input perturbations comparing single-variable, one-hot vector, and population-code outputs. Here, the networks were trained by adding Gaussian input noise. Experimental results show means $\pm$ SD indicated by shaded areas (100 simulation runs).
@@ 3.3 Population Code vs One-Hot Vector
So, we suspect that the one-hot method has larger linear-layer outputs compared to the population code. The larger values cause a problem for robustness because they make it more likely that a perturbation causes spurious activations that suppress the correct output value. To test this hypothesize, we evaluated the activations after the final linear layer, and compared one hot with population code. As a result, both the max and min values were indeed larger for one hot for the 3, 5, and 8-layer networks (Fig. 4). For the 8-layer network, the values were about 5x larger.
@@ 4.3 Experimental Results
Table 2: Comparison of methods for object 5 in the T-LESS dataset. For the single-variable approach, we used multiple target to deal with the pose ambiguity.

### v4

@@ 1 Introduction
In the mammalian cortex, many variables, e.g., object orientation [13] and movement direction [6], have been found to be encoded by populations of neurons. In a population code, each neuron responds maximally to its preferred value of an encoded variable and partially to other values. Activation levels are shaped by tuning curves such as Gaussian or cosine functions of the distance between encoded and preferred value. The activity of the group of neurons resembles a probability distribution of the encoded variable [18].
@@
Commonly used artificial neural networks, such as convolutional neural networks (CNNs) and multi-layer perceptrons (MLPs), encode information also by groups of neurons, particularly in their intermediate layers [8]. However, in the output layer, information is typically mapped onto a different representation. For classification tasks, outputs are commonly represented as one-hot vectors, using one neuron for each classification label. For prediction tasks, output neurons typically correspond to the variables of interest, such as the position and orientation of an object in an image.
@@ 2.1 Single-Variable Output
For the weights and bias to converge, the following equality must hold,
@@
Since this equality can be fulfilled for any $b$ value, the weights are defined only up to a constant bias term that shifts all weight values. The actual value of $b$ will depend on network initialization.
Next, we probe the network’s robustness to noise in the input. For this test, we perturb the input ${\bf x}$ by ${\bf d}$, so that ${\bf x}+{\bf d}$ is the new input. Moreover, let $d_{j}=a\delta_{jk}$ for a given perturbed neuron $k$, and we use the same ${\bf x}$ and target, $i/n$, as above. As a result, the output equals,
@@ 2.2 Population-Code Output
$w_{oi}^{t+1}=w_{oi}^{t}+\eta\left(\hat{y}_{o}-\sum_{j}w_{oj}^{t}x_{j}-b_{o}\right)x_{i}\,.$ (12)
For our theoretical analysis, we first use a simplified population code that resembles a one-hot vector, i.e., only one neuron, $y_{o}$, with preferred value $o/n$ gets active, and all other neuronal activations are zero. Later, we will show the difference between population code and one-hot vector.
@@
Using the training input, $x_{j}=\delta_{ij}$ and target $\hat{y}_{o}=\delta_{oi}$ with input value 1 at index $i$, the weight update simplifies to
$w_{oi}^{t+1}=w_{oi}^{t}+\eta\left(\delta_{oi}-w_{oi}^{t}-b_{o}\right)\,.$ (13)
@@
As above, we perturb the input by $d_{j}=a\delta_{jk}$. Using (11) and (14), the perturbed input produces the output
$y_{o}=\delta_{oi}+a\delta_{ok}-ab_{o}\,.$ (16)
For a failure to occur, a necessary condition is $y_{k}>y_{i}$. Using (16), this condition is equivalent to
$a>\frac{1}{1-b_{k}+b_{i}}$ (17)
if $01-b_{k}+b_{i}>$ 0. Since the bias is commonly initialized to zero and due to the symmetry of our problem with respect to the indices, we expect only a minor difference between bias values. Therefore, $a$ has to be near 1 for a failure to occur, which makes the population code more robust to noise compared to the single variable output.
For an actual population code, the activation values follow a tuning curve rather than being binary. So, the allowable threshold for $a$ will be slightly smaller, as shown in the following. For Gaussian tuning, a failure occurs if the maximum of the sum of the Gaussians from the real signal and perturbation is shifted by at least $2/n$ compared to the target value (the preferred values are discrete in steps of $1/n$). This condition necessitates that the activation at target is smaller than this maximum,
@@
where $\sigma$ is the tuning width, and $4$ is the worst-case difference between $i$ and $k$ that could result in errors above the tolerance of $1.5/n$ for $a<1$. The right-hand side is maximized at $x=2/n$ because the maximum cannot be at $x>2/n$ if $a<1$ due to the higher amplitude of the Gaussian at target ($x=0$).
@@ 3.1 Methods
We used various numbers of stacked linear layers of size $n=20$, except for the output layer for the single-variable approach, which consisted of only one neuron. In the architecture, each hidden layer was followed by a leaky ReLU function. For the population-code approach, a sigmoid function was applied to the output, and we used Gaussian tuning curves with a width of $\sigma=0.1$.
@@
Figure 1: Robustness to input perturbations comparing single-variable, one-hot vector, and population-code outputs. Here, the networks were trained without input noise. The dashed thin red line in the first panel shows the theoretical bound of $a$ below which the population-code has zero errors. Experimental results show means $\pm$ SD, as indicated by shaded areas (100 simulation runs).
@@ 3.2 Results on Noise Robustness
Figure 2: Robustness to input perturbations comparing single-variable, one-hot vector, and population-code outputs. Here, the networks were trained by adding Gaussian input noise. Experimental results show means $\pm$ SD, as indicated by shaded areas (100 simulation runs).
@@ 3.3 Population Code vs One-Hot Vector
So, we suspect that the one-hot method has larger linear-layer outputs compared to the population code. The larger values cause a problem for robustness because they make it more likely that a perturbation causes spurious activations that suppress the correct output value. To test this hypothesize, we evaluated the activations after the final linear layer, and compared one hot with population code. As a result, both the absolute max and min values were indeed larger for one hot for the 3, 5, and 8-layer networks (Fig. 4). For the 8-layer network, the values were about 5x larger.
@@ 4.3 Experimental Results
Table 2: Comparison of methods for object 5 in the T-LESS dataset. For the single-variable approach, we used multiple targets to deal with the pose ambiguity.